# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v1

@@ VI-A Randomized Maximal Inscribed Ellipse Algorithm via Bottom-Up Strategy
5      $\mathcal{S}\leftarrow\{\}$ $\mathcal{B}\leftarrow\mathcal{X}$ $v_{\mathcal{B}}\leftarrow v_{\mathcal{X}}$ foreach $h\in\mathcal{H}~{}~{}\mathrm{in~{}a~{}random~{}order}$ do
@@
7                   $\mathcal{B},v_{\mathcal{B}}\leftarrow\mathbf{MaxEllipse}(\mathcal{S},\mathcal{X}\cup\{h\},v_{\mathcal{X}})$
@@ 9            $\mathcal{S}\leftarrow\mathcal{S}\cup\{h\}$
Input: $\mathcal{X},v_{\mathcal{X}}$
1 begin
2       if $\mathbf{CheckCloesd}(\mathcal{X})$ then
3             $\mathbf{MENN}(|\mathcal{X}|,\mathcal{X},v_{\mathcal{X}})$    $\triangleright$ detailed in Sec.VI-B
4      else
5             $\mathbf{ConeAngle}(\mathcal{X},v_{\mathcal{X}})$
6       end if
7
8 end
9
Algorithm 4 UpdateValue
Then we discusse the primitive operations involved in the above problem by case. For the case when $\mathcal{H}$ is unclosed, both primitives are trivial. For the closed case, the violation test can be implemented by evaluating (30b), but the basis computation is challenging. The basis here are actually a subset of the edges tangent to the MVIE of $\mathcal{H}$. In the original MSW algorithm, the basis of $\mathcal{H}$ will be directly computed by (VI-A-47), which involves huge amounts of redundant calculations. Thus, we here utilize the bottom-up strategy aforementioned to rebuild the recursive algorithm of MSW, utilizing the property that the basis computation of each subset of $\mathcal{H}$ is part of basis computation of $\mathcal{H}$. Finally, a novel randomized algorithm we proposed for 2-D MVIE are presented in Algorithm 3 and 4.
Algorithm 3 takes as input the set $\mathcal{H}$ to be checked, the set $\mathcal{X}$ known to belong to the basis of $\mathcal{H}\cup\mathcal{X}$, and $v_{\mathcal{X}}$ that will update at the beginning of the algorithm (line 3), which is detailed in Algorithm 4. The output of Algorithm 3 is the basis of $\mathcal{H}\cup\mathcal{X}$ and its value. For the initial call to this algorithm, we set $\mathcal{H}\leftarrow\mathcal{P}$ and $\mathcal{X}\leftarrow\{\}$. Within the loop (line 3-3), each elements $h$ of $\mathcal{H}$ is checked sequentially in a randomized order, testing whether $h$ is violated by $\mathcal{B}$. If violated, we recursively call the algorithm (line 3). Based on the enumeration principle in Fig. 3, the number of the elements of $\mathcal{X}$ in the recursive algorithm cannot exceed the combination dimension $\delta$ (line 3). In ViolateTest function (line 3), in addition to the evaluation involved $h$ and $v_{\mathcal{B}}$ based on (48-51), in the case where $h$ is violated by $\mathcal{B}$, if $\mathcal{X}\cup\{h\}$ forms a convex $N$-gon but $N<|\mathcal{X}\cup\{h\}|$, then false is returned. The reason is that we perform checks based on the bottom-up strategy, the non-redundant closed subset of the closed set $\mathcal{X}\cup\{h\}$ that contains redundant elements will definitely be checked in the algorithm. Therefore, there is no need to check on the redundant closed set. Moreover, this ensures that the input $\mathcal{X}$ of Algorithm 4 is free of redundant elements in the closed case. Overall, the thought of bottom-up is embedded in the violation test and recursive calls. For example, when we recursively call Algorithm 3 with the input that meets $|\mathcal{X}|=4$, it means that the basis of each its subset containing $3$ edges has already been checked and violated by the fourth constraint. Thus, in Algorithm 4, if the input $\mathcal{X}$ is closed, we can directly compute its MENN as its MVIE to update $v_{\mathcal{X}}$. Otherwise we will compute the angle value of $\mathcal{X}$ to update $v_{\mathcal{X}}$.
By introducing the bottom-up strategy and reformulating MVIE into an LP-type problem, we avoid the need for violent enumeration and redundant calculations in solving (VI-A-47). Eventually, we achieve linear-time complexity for solving 2-D MVIE. Another factor that affects the efficiency of the algorithm is the MENN subproblem (VI-A) solving, for which we present an efficient analytic solution in Sec. VI-B.
VI-B Maximal Ellpise tangent to N edges of the N-gon
@@
As demanded in Algorithm 4, in this section we present the analytic computation of the maximal inscribed ellipse that is tangent to all edges of arbitrary convex $N$-gon, with $N\in\{3,4,5\}$, which we denote as MENN. For arbitrary non-degenerate triangle [41], convex quadrilateral [42] or convex pentagon [35], there exits and only exits unique such ellipse. Since the inputs from Algorithm.4 are non-degenerate and convex, in the following we default to the existence and uniqueness of the MENN.
@@
Referring to the definition in (4), in the 2-D case we define the point $p$ on the boundary of the ellipse $\mathcal{E}$ to satisfy
@@
We denote the homogeneous coordinate [43] of point $p$ as
$\hat{p}=\left(p_{0},~{}p_{1},~{}1\right)^{\mathrm{T}},$ (53)
where *$*_{i}$ represents the $i$-th element of the vector ***. Since $A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}$ is positive definite symmetric, (52) is a second-degree polynomial equation in two variables $p_{0}$ and $p_{1}$. We transform (52) into a homogeneous quadratic form, then the ellipse $\mathcal{E}$ can be described in homogeneous coordinate as
=$\mathbb{P}=\left\{{\hat{p}~{}\big{|}~{}\hat{p}^{\mathrm{T}}M_{P}\hat{p}=0}\right\},$ (54)
@@
based on which, the calculation of the line coordinate of a line can be performed by the Grassmanian expansion of the homogeneous cooridinates of two different points on the line [45]. Then, according to the polarity of ellipse [44], when $\hat{p}$ fulfils (54), which means $\hat{p}$ is on the ellipse $\mathcal{E}$, the line coordinate $l$ of the line tangent to $\mathcal{E}$ at the point $\hat{p}$ is given as
@@
$\mathbb{L}=\left\{{l~{}\big{|}~{}l^{\mathrm{T}}M_{L}l=0}\right\},$ (57)
@@
where the coefficient matrix $M_{L}\in\mathbb{R}^{3\times 3}$ is symmetric.
Now we can calculate the MENN by utilizing the information of the tangents directly, abandoning the necessity of calculating the points of tangency. Specifically, we first compute the line coordinates of edges of the polygon, which actually are the tangents of the ellipse, through the vertices of the polygon by (55). Then we calculate the coefficient matrix $M_{L}$ via the tangents in the line coordinate by (57), based on which the coefficient matrix $M_{P}$ in the homogeneous corrdinate can be obtained by (58). Finally the ellipse in the desired form of (4) can be calculated by (52) and (54). In the following, we present the implementation details.
VI-B1 MENN of a convex pentagon
For (57), the symmetric $M_{L}$ has $6$ variables. Bringing the line coordinates of the five edges of the convex pentagon into (57) respectively, we can obtain a six-element homogeneous system of linear equations consisting of five equations. Since there exits and only exits unique ellipse tangent to all edges of the convex pentagon [35], the homogeneous system always has a non-trivial solution. Then based on the solution, we can obtain the MENN progressively through the process aforementioned.
VI-B2 MENN of a convex quadrilateral
In contact to convex pentagon, a covex quadrilateral only provide four tangents. Thus there are a unique one-parameter set of inscribed ellipses that are tangent to all edges of the quadrilateral, whose parameter can be taken to be a prescribed point of contact on any single edge of the quadrilateral [35] as shown in Fig. 5. Additionally, only one of them has the largest area [42], which is the MENN of the quadrilateral.
Figure 5: Illustration of the calculation of the MENN of a convex quadrilateral. The blue and red ellipses represent the ellipses obtained by sampling different points $\hat{p}_{\lambda}$ on the coinciding edge $l_{\lambda}$.
@@ VI-B2 MENN of a convex quadrilateral
$\hat{p}_{\lambda}=(\lambda,0,1)^{\mathrm{T}},~{}~{}l_{\lambda}=[0,1,0]^{\mathrm{T}},$ (59)
where $\lambda$ is a variable indicating the position of the point. Combining (56) and (58), their polarity relationship can be written as
@@
which is an independent new linear constraint on variables of $M_{L}$. By incorporating this constraint with the equations get by substituting the four tangents into (57), we acquire a system of equations similar to Sec. VI-B1. Via the previously described process, we obtain $M_{L}$ and $M_{P}$ sequentially which here are in trems of $\lambda$. Based on $M_{P}$, the area $A_{\mathcal{E}}$ of ellipse $\mathcal{E}$ [46] can be calculated by
$A_{\mathcal{E}}(\lambda)=\frac{\det({M_{P}})}{{\det({\tilde{M}_{P}})^{\frac{3}{2}}}}\pi,$ (61)
where $\tilde{M}_{P}\in\mathbb{R}^{2\times 2}$ is the submatrix obtained by removing the leftmost column and the top row of $M_{P}$. Finally, the optimal $\lambda$ corresponding to the MENN of the convex quadrilateral can be computed by calculating the zeros of the first order derivative of $A_{\mathcal{E}}(\lambda)$. This calculation only involves solving a quadratic equation w.r.t. $\lambda$, which can be solved quickly and analytically.
VI-B3 MENN of a triangle
For a triangle, the Steiner inellipse which is tangent to the three edges of the triangle at their midpoints, has the largest aera [41], which is the MENN of the triangle. We denote the 2-D Cartesian coordinates of the vertices of the triangle as $v^{*}=(v^{*}_{0},v^{*}_{1})^{\mathrm{T}},*=\{\textrm{i},\textrm{ii},\textrm{iii}\}$. The center $v^{o}$ and two conjugate diameters $f_{1},f_{2}\in\mathbb{R}^{2}$ of the Steiner inellipse can be written as
@@ VI-B3 MENN of a triangle
$b_{\mathcal{E}}=v^{o},~{}A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}=\bigl{(}f_{1}~{}f_{2}\bigr{)}\bigl{(}f_{1}~{}f_{2}\bigr{)}^{\mathrm{T}}$ (64)
VII Evaluation and Benchmark
In order to comprehensively prove the superiority of the algorithms proposed in this paper, we design comparison experiments for solving MVIE and for generating convex free polytope. We will demonstrate the efficiency of the proposed methods in the benchmark of solving MVIE, which greatly benefits the computational efficiency of the proposed algorithm for convex free space generation, FIRI. Then we will show the outstanding advantages of FIRI in terms of computational efficiency, quality and managibility by comparing FIRI with other state of the art convex polytope generation algorithms extensively. All benchmarks are conducted on an Intel(R) Core(TM) i7-12700KF CPU @ 3.61GHz under Linux, and are implemented in C++11 without any explicit hardware acceleration.
Figure 6: Computation time of different methods for MVIE under different constraint numbers.
TABLE I: Comparison of Precision $\psi_{\mathcal{E}}$ between Different Methods for Solving MVIE
Scenario Precisions $\psi_{\mathcal{E}}$
IRIS[1] Mosek[19] SDP Mosek[19] SOCP CAS (Ours) RAN (Ours)
2-D 8.56e-8 6.47e-9 4.87e-12 1.59e-8 4.41e-16
3-D 1.54e-8 1.56e-8 4.05e-12 2.04e-8 /
VII-A Comparison of Solving MIVE
For MVIE (11), we propose two algorithms, namely the SOCP-reformulation algorithm and the randomized algorithm specialized for 2-D case, which are presented in Sec. V and VI, respectively. Since the randomized algorithm yields an analytical solution, here we abbreviate it as RAN. Then to distinguish from the methods to be compared, we denote the proposed SOCP-reformulation algorithm which is solved by affine scaling method as CAS. We compare the proposed algorithms with three methods: 1) The optimization method based on SDP formulation of MVIE in IRIS [1]. 2) The example  [^1] https://docs.mosek.com/latest/cxxfusion/examples-list.html#doc-example-file-lownerjohn-ellipsoid-cc of the cutting-edge solver Mosek [19] to computes the Lowner-John inner ellipsoidal approximations of a polytope. 3) A strategy of solving the SOCP form (V) of the MVIE by using Mosek. As the example in Mosek formulate MVIE into a mixed conic quadratic and semidefinite problem, we denote it as Mosek SDP below and denote another strategy using Mosek as Mosek SOCP. We use the default parameters for IRIS and Mosek.
where ${A_{\mathcal{E}}}^{*}$, ${D_{\mathcal{E}}}^{*}$ and ${b_{\mathcal{E}}}^{*}$ are the coefficients of the maximum inscribed ellipsoid solved by each method, $A_{\mathcal{P}}$ and $b_{\mathcal{P}}$ are the parameters that define the halfspaces of the input polytope $\mathcal{P}$ (10), and $\mathcal{L}_{\psi}(\cdot)$ denotes a function that takes the absolute value of the largest element of the input vector. The corresponding performance are summarized in Fig. 6 and Tab. I. By comparing the results of Mosek SDP and Mosek SOCP, we can get this conclusion that transforming the commonly used SDP formulation to the SOCP formulation leads to a significant improvement in computational efficiency without sacrificing accuracy. Additionally, in the low-dimension massive-constraint case faced in this paper, the use of the affine scaling method avoids the requirement of solving a large-scale system of linear equations at each iteration, compared to the interior point method used in Mosek. Thus as shown the results demonstrate, CAS is capable of solving MVIE in less time while maintaining a comparable precision. Moreover, for the 2-D case, our proposed linear-time complexity algorithm RAN further enhances the computational efficiency by orders of magnitude and compute analytical results.
VII-B Comparison of Generating Convex Free Polytope
Based on the description and analysis of several state-of-the-art algorithms for convex free polytope generation in Sec. II, we benchmark the proposed algorithm FIRI whith IRIS [1], Galaxy [3] and RILS [4]. Since Gao’s method [2] relies on modeling the environment as a grid map and can only be barely real-time with low map resolution, it is not compared here. Additionally, Galaxy can be regarded as an enhancement of Savin’s method [8], both of whcih are based on space inversion, hence we directly employ Galaxy. For IRIS and our proposed method FIRI, we set the same parameter $\rho=0.02$, that is, the stopping condition of them is that the volume of the MVIE of the convex hull obtained in this iteration grows less than $2\%$ from the last iteration. For Galaxy and RILS, we used the default parameters.
TABLE II: Comparison of Adaptability of Different Methods for Generating Convex Free Polytope
Method Adaptability Seed Type Obstacles Type
  Point Line Polytope Point Polytope
FIRI
IRIS[1]
Galaxy[3]
RILS[4]
TABLE III: Success Rate of Different Methods for Generating Convex Polytopes Containing Seed across Different Scenarios and Seed Types

### v2

@@ 2       $v_{\mathcal{X}}\leftarrow w(\mathcal{X})$   /* defined in (44) */
6      $\mathcal{S}\leftarrow\{\}$
7       $\mathcal{B}\leftarrow\mathcal{X}$
8       $v_{\mathcal{B}}\leftarrow v_{\mathcal{X}}$
9       foreach $h\in\mathcal{H}~{}~{}\mathrm{in~{}a~{}random~{}order}$ do
@@ 8       $v_{\mathcal{B}}\leftarrow v_{\mathcal{X}}$
11                   $\mathcal{B},v_{\mathcal{B}}\leftarrow\textbf{MaxEllipse}(\mathcal{S},\mathcal{X}\cup\{h\})$
12
@@ 14            $\mathcal{S}\leftarrow\mathcal{S}\cup\{h\}$
VI-B3 Algorithm Overview
Building upon the aforementioned efforts, we improve the general framework [26] to propose a randomized algorithm for solving 2-D MVIE as shown in Algorithm 3. For the initial call to Algorithm 3, we set $\mathcal{H}\leftarrow\mathcal{P}$ and $\mathcal{X}\leftarrow\{\}$. The recursive framework (Line 3) of the algorithm, combined with the previously mentioned bottom-up strategy, results in an effect: when a recursive call occurs, it indicates that all subsets of the input $\mathcal{X}$ have already been checked, and for any of its subsets $\bar{\mathcal{X}}$, there is always an element $h\in\mathcal{X}$ violated by $\bar{\mathcal{X}}$. Therefore, if the input $\mathcal{X}$ is closed, we can directly solve the MENN of the polygon formed by $\mathcal{X}$ for the value $w(\mathcal{X})$ in Line 3. And if $\mathcal{X}$ is unclosed, it is easy to use (45) to compute its value. Recalling the complexity conclusion we mentioned in Sec. VI-A, now we can obtain the result of Algorithm 3 by finite number of violation tests (Line 3), MENN (for closed case) and (45) (for unclosed case), whose expected number is linear in the input set number.
In addition, the maximum cardinality of any basis is denoted as combinatorial dimension $\delta$, and based on the value function (44) we have $\delta=5$. This implies that in Line 3, we will only encounter MENN calculations for triangles, quadrilaterals, and pentagons, which also consistent with our decomposition of MVIE in (48). The combinatorial dimension $\delta$ in Line 3 ensures that the maximum number of recursive iterations in the algorithm remains a constant.
The analytical solution method (detailed in Sec. VI-C) may lead to incorrect result when the input is redundant. Thus, as mentioned in Sec. VI-B2, by treating redundant cases separately and assigning them a solution of $\infty$, we prevent them from influencing the overall algorithm. In Algorithm 3, we add this feature in ViolateTest (Line 3): in addition to the evaluation involved $h$ and $v_{\mathcal{B}}$ based on (44), in the case where $h$ is violated by $\mathcal{B}$, if $\mathcal{X}\cup\{h\}$ forms a convex $N$-gon but $N<|\mathcal{X}\cup\{h\}|$ (redundant case), then false is returned. Since we perform checks based on the bottom-up strategy, the non-redundant closed subsets (if they exist) of the redundant closed set $\mathcal{X}\cup\{h\}$ will definitely be checked in the algorithm. Thus there is no need to check with the redundant closed set.
Now we construct the linear complexity algorithm for 2-D MVIE. Another factor that affects the practical efficiency of the algorithm is the solution of MENN (48), for which we present efficient analytical solutions in Sec. VI-C.
VI-C Maximal Ellpise tangent to N edges of the N-gon
@@
As demanded in Algorithm 3, in this section we present the analytical computation of MENN, which is the maximal inscribed ellipse that is tangent to all edges of arbitrary convex $N$-gon, $N\in\{3,4,5\}$. For arbitrary non-degenerate triangle [49], convex quadrilateral [50] or convex pentagon [42], there exists a unique such ellipse. Since the inputs from Algorithm 3 are non-degenerate and convex, in the following we default to the existence and uniqueness of the MENN.
@@
Referring to the definition in (4), in the 2-D case we define the point $p\in\mathbb{R}^{2}$ on the boundary of the ellipse $\mathcal{E}$ to satisfy
@@
We denote the homogeneous coordinate [51] of point $p$ as $\hat{p}=\left(p^{\mathrm{T}},~{}1\right)^{\mathrm{T}}$. Since $A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}$ is positive definite symmetric, (50) is a second-degree polynomial equation in two elements of $p$. We transform (50) into a quadratic form, then the ellipse $\mathcal{E}$ can be described in homogeneous coordinate as
=$\mathbb{P}=\left\{{\hat{p}\in\mathbb{R}^{3}~{}\big{|}~{}\hat{p}^{\mathrm{T}}M_{P}\hat{p}=0}\right\},$ (51)
@@
based on which, the calculation of the line coordinate can be performed by the Grassmanian expansion [53]. According to the polarity of ellipse [52], when $\hat{p}$ is on the ellipse $\mathcal{E}$ (51), the line coordinate $l$ of the line tangent to $\mathcal{E}$ at $\hat{p}$ is given by
@@
$\mathbb{L}=\left\{{l\in\mathbb{R}^{3}~{}\big{|}~{}l^{\mathrm{T}}M_{L}l=0}\right\},$ (54)
@@
Now we can calculate the MENN by utilizing the information of the tangents directly. Specifically, we first compute the line coordinates of edges of the polygon for solving $M_{L}$, then obtain $M_{P}$, and eventually get the ellipse in the desired form of (4). This process involves using (52), (54), (55), (50) and (51) sequentially. In the following, we present the implementation details for different $N$.
VI-C1 MENN of a Convex Pentagon
For (54), the symmetric $M_{L}$ has $6$ variables. Bringing the line coordinates of the five edges of the convex pentagon into (54) respectively, we can obtain a six-element homogeneous system of linear equations consisting of five equations. Since there exists and only exists unique ellipse tangent to all edges of the convex pentagon [42], the homogeneous system always has a non-trivial solution. Then based on the solution, we can obtain the MENN progressively through the process aforementioned.
VI-C2 MENN of a Convex Quadrilateral
In contrast to convex pentagon, a convex quadrilateral only provides four tangents. Thus there are a unique one-parameter set of inscribed ellipses that are tangent to all edges of the quadrilateral, whose parameter can be taken to be a prescribed point contact on any single edge of the quadrilateral [42] as shown in Fig. 5. Additionally, only one of them has the largest area [50], which is the MENN of the quadrilateral.
Figure 5: Left: Illustration of the calculation of the MENN of a convex quadrilateral. The blue and red ellipses represent the ellipses obtained by sampling different points $\hat{p}_{\lambda}$ on the coinciding edge $l_{\lambda}$. Right: Illustration of the calculation of the MENN of a triangle.
@@ VI-C2 MENN of a Convex Quadrilateral
$\hat{p}_{\lambda}=(\lambda,0,1)^{\mathrm{T}},~{}~{}l_{\lambda}=(0,1,0)^{\mathrm{T}},$ (56)
where $\lambda$ is a variable. Combining (53) and (55), their polarity relationship can be written as
@@
which is an independent new constraint on $M_{L}$. The subsequent operations are similar to Sec. VI-C1, with the difference that eventually we calculate the area [54] of $\mathcal{E}$ as a function $A_{\mathcal{E}}(\lambda)$ in terms of $\lambda$. The optimal $\lambda$ corresponding to the MENN of the quadrilateral can be computed by calculating the zeros of the first order derivative of $A_{\mathcal{E}}(\lambda)$. This calculation only involves solving a quadratic equation w.r.t. $\lambda$, which can be solved quickly and analytically. Due to space limitations, we do not delve into further details of the calculation.
VI-C3 MENN of a Triangle
The Steiner inellipse which is tangent to the edges of the triangle at their midpoints, is the MENN of the triangle [49]. As shown in Fig. 5, we denote the 2-D Cartesian coordinates of the vertices of the triangle as $v^{*}=(v^{*}_{0},v^{*}_{1})^{\mathrm{T}},*=\{\textrm{i},\textrm{ii},\textrm{iii}\}$. The center $v^{o}$ and two conjugate diameters $f_{1},f_{2}\in\mathbb{R}^{2}$ of the Steiner inellipse can be written as
@@ VI-C3 MENN of a Triangle
$b_{\mathcal{E}}=v^{o},~{}A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}=\bigl{(}f_{1}~{}f_{2}\bigr{)}\bigl{(}f_{1}~{}f_{2}\bigr{)}^{\mathrm{T}}.$ (60)
TABLE I: Success Rate of Different Methods for Generating Convex Polytopes Containing Seed across Different Scenarios and Seed Types
@@
VII Evaluation and Benchmark
To comprehensively demonstrate the outstanding advantages of the proposed algorithm, FIRI, in terms of efficiency, quality, and manageability, we design extensive benchmarks comparing FIRI against several state-of-the-art convex polytope generation algorithms. Additionally, the efficiency of FIRI lies in the methods that we develop specifically for the two optimization problems (Line 1 and 1 of Algorithm 1) within FIRI. To demonstrate their effectiveness, we conduct comparative experiments for these methods which are proposed for solving minimum-norm with small dimension but massive constraints and for solving MVIE, respectively. All benchmarks are conducted on an Intel Core i7-12700KF CPU under Linux, and are implemented in C++11 without any explicit hardware acceleration.
VII-A Comparison of Generating Free Convex Polytope
Based on the description and analysis of several state-of-the-art algorithms for generating free convex polytopes in Sec. II, we benchmark the proposed algorithm FIRI against IRIS [1], Galaxy [9] and RILS [2]. Since Gao’s method [8] relies on modeling the environment as a grid map and can only operate in near real-time with low map resolution, it is not compared here. Additionally, Galaxy can be regarded as an enhancement of Savin’s method [12], both of which are based on space inversion, thus we directly adopt Galaxy for benchmark. For IRIS and our proposed method FIRI, we set the same parameter $\rho=0.02$, that is, the stopping condition of them is that the volume of the MVIE of the convex hull obtained in this iteration grows less than $2\%$ from the last iteration. For Galaxy and RILS, we used the default parameters.